OpenStudy (anonymous):

What is the missing integer in this problem? ____- 8 = -5

OpenStudy (anonymous):

This is essentially the same as x - 8 = -5 So, add 8 to both sides. x - 8 + 8 = -5 + 8 x = 3

OpenStudy (anonymous):

thats really all i had to do????

OpenStudy (anonymous):

Yep! You're looking for an unknown, (x), and when subtracted by 8, is equal to -5. Just treat it like a linear equation!

OpenStudy (anonymous):

easier way ... -5+8= 3

OpenStudy (anonymous):

thank you !!!!! xo

OpenStudy (anonymous):

You're welcome!